The crown of a straw hat has a base of 38 square inch.The depth of the crown is 3 inches.If the head occupies two thirds of the space enclosed by the crown,find the volume remaining for ventilation.
Answer by checkley77(12844)

38*3=114 cubic inches.
114*2/3=228/3=76 cubic inches for the head space.